# Certified Causal Defense with Generalizable Robustness

### v2

@@ Abstract
While machine learning models have proven effective across various scenarios, it is widely acknowledged that many models are vulnerable to adversarial attacks. Recently, numerous efforts have emerged in adversarial defense. Among them, certified defense is well known for its theoretical guarantees against arbitrary adversarial perturbations on input within a certain range. However, most existing works in this line struggle to generalize their certified robustness in other data domains with distribution shifts. This issue is rooted in the difficulty of eliminating the negative impact of spurious correlations on robustness in different domains. To address this problem, in this work, we propose a novel certified defense framework GLEAN, which incorporates a causal perspective into the generalization problem in certified defense. More specifically, our framework integrates a certifiable causal factor learning component to disentangle the causal relations and spurious correlations between input and label, thereby excluding the negative effect of spurious correlations on defense. On top of that, we design a causally certified defense strategy to handle adversarial attacks on latent causal factors. In this way, our framework is not only robust against malicious noises on data in the training distribution but also can generalize its robustness across domains with distribution shifts. Extensive experiments on benchmark datasets validate the superiority of our framework in certified robustness generalization in different data domains.
@@ Introduction
Machine learning (ML) models, particularly deep neural networks (DNN), have demonstrated remarkable success across a diverse range of areas (Devlin et al. 2018; Silver et al. 2017; He et al. 2016). Despite their success, these models still exhibit significant vulnerabilities to adversarial perturbations on input (Szegedy et al. 2013; Goodfellow, Shlens, and Szegedy 2014; Biggio et al. 2013). A typical example in image classification is that a trained classifier that correctly classifies an image $x$ can be easily fooled by a perturbed image $x+\delta$, where $\delta$ represents adversarial perturbations that are imperceptible to human perception. This weakness impedes the deployment of ML models in critical applications where security and reliability are priorities, such as autonomous driving and healthcare.
@@
In the past few decades, researchers have developed numerous defense methods to enhance the adversarial robustness of ML models. Many of these methods are based on adversarial training (Goodfellow, Shlens, and Szegedy 2014; Madry et al. 2017; Zhang et al. 2019; Athalye, Carlini, and Wagner 2018), which incorporates adversarial samples into model training. Despite its impressive performance, adversarial training is an empirical approach that lacks theoretical guarantees. That is, although it can enhance robustness against certain types of attacks, it may still be vulnerable to other unknown, or more potent adversarial perturbations. Differently, another line of work develops certified robustness. A certified robust classifier can theoretically guarantee that its prediction for a point $x$ remains constant within a certain specified range (a.k.a. radius) of perturbations on $x$, regardless of the type of attack. Randomized smoothing-based certified defense (Lecuyer et al. 2019; Li et al. 2018; Cohen, Rosenfeld, and Kolter 2019) is one of the most representative methods in this area. Specifically, given an arbitrary base classifier $f$, this method can convert it to a certifiably robust classifier $g$, which is created by randomly sampling multiple noised versions of a given input and using the aggregated output from these variations to make final predictions. Inspired by this approach, many subsequent studies (Li et al. 2019; Jeong and Shin 2020; Jeong et al. 2021; Salman et al. 2019; Zhai et al. 2020) have expanded upon the basis of random smoothing.
@@
Currently, most existing certified defense works focus on data in the same domain yet overlook other domains with distribution shifts. This limitation can result in a markedly degraded certified robustness performance when these methods are applied to the test domain (Sun et al. 2021). As discussed in previous work (Ilyas et al. 2019; Beery, Van Horn, and Perona 2018), such degradation of robustness lies in the fact that ML models tend to overfit spurious correlations between features and labels. As these spurious correlations often vary across different domains (Ye et al. 2022), fitting spurious correlations can easily lead $g$ to make incorrect predictions or correct predictions but with lower confidence levels. The former results in the certified radius being assigned zero, while the latter also leads to a reduced certified radius. Therefore, domain shifts can lead to weak generalization w.r.t. not only prediction performance but also certified robustness. Different from ML models, humans can naturally capture the invariant relations between labels and their corresponding causal factors, various studies (Zhang, Zhang, and Li 2020; Tang, Huang, and Zhang 2020; Schölkopf et al. 2021) argue that human’s inherent causal view brings a solution to avoid the domain generalization hurdle for robustness. Inspired by this, in this paper, we study the problem of generalizing the certified robustness under domain shifts from a causal view.
@@
However, addressing this problem presents multifaceted challenges. Challenge 1: As aforementioned, spurious correlations varying across domains adversely affect robustness. To achieve robustness across domains, it is crucial to effectively remove the impact of these spurious correlations. However, identifying and eliminating the impact of spurious correlations on robustness in different domains presents a challenge. Challenge 2: Apart from spurious factors, the distribution shifts often bring challenges for the model to defend against the perturbations on the factors that causally determine the label in unseen domains, which leads to diminished certified robustness. Challenge 3: It is important to provide theoretical guarantees for robustness on other data domains, but most existing works remain in empirical observations and lack theoretical analysis. Although some works (Salman et al. 2019; Zhai et al. 2020) in certified defenses has provided upper bounds on perturbations while maintaining robustness, they were not designed to address certified robustness in the domain shift context.
@@
In this work, to tackle these challenges, we propose a novel framework GeneraLizable cErtified cAusal defeNse (GLEAN) that enhances the certified robustness of models on data in different domains. To mitigate the influence of spurious correlations on robustness generalization (Challenge 1), we construct a causal model for data in different domains, and simultaneously conduct a causal analysis on model robustness and generalization. Based on the causal model, we filter out the impact from spurious correlations and enhance robustness across domains. This is different from most existing defense algorithms which take the same strategy indiscriminately towards all the input features. To achieve certified robustness through causal factors (Challenge 2), we utilize a certified causal factor learning module with Lipschitz constraint. This module enables certification through the latent representation space for high-level causal factors, conducting certified defense for perturbations on causal factors that determine the label in different domains. To bring theoretical guarantees for robustness on different data domains (Challenge 3), we derive a theoretical analysis by leveraging the theoretical support of certified defense and causal inference. Our main contributions can be summarized as:
@@
Many previous works proposed certification methods to derive a certified radius that is the lower bound of the robust radius. Research in this area falls into two categories: exact methods and conservative methods. Exact methods (Ehlers 2017; Bunel et al. 2018; Tjeng, Xiao, and Tedrake 2017), usually based on Satisfiability Modulo Theories or mixed integer linear programming, guarantee the identification of a perturbation $\delta$ within a radius $r$ that can cause $f$ to change its prediction. However, they require the model to have a limited scale. Conservative methods (Wong and Kolter 2018; Wong et al. 2018; Gowal et al. 2018) ensure the detection of existing adversarial examples and, in addition, refuse to make certification for some vulnerable data points. These methods, though more scalable, impose specific assumptions on the model’s architecture.
@@
Randomized Smoothing (RS) is proposed to tackle the above limitations, which can be applied to any architectures (Cohen, Rosenfeld, and Kolter 2019). It constructs a smoothed classifier $g$ from an arbitrary base classifier $f$. The definition of $g$ is as follows:
@@
In this formula, $\eta\sim\mathcal{N}(0,\sigma^{2}\bm{I})$ is the isotropic Gaussian noise with the noise level $\sigma$ as the hyperparameter. The smoothed classifier $g$ can be summarized as returning the class most likely to be predicted by $f$ when the input $x$ is sampled from Gaussian distributions. (Cohen, Rosenfeld, and Kolter 2019) provide the theoretical form of certified radius which is the lower bound of the robust radius:
@@
where $p_{A}=P(f(x+\eta)=y_{A})$, $p_{B}=\max_{y\neq y_{A}}P(f(x+\eta)=y)$, meaning that $f$ will mostly return class $y_{A}$ with the probability $p_{A}$, and will return the runner-up class with the probability $p_{B}$. $\Phi^{-1}$ is the inverse of the standard Gaussian cumulative distribution function. Then $g(x+\delta)=y_{A}$ for all $||\delta||_{2}\leq{CR}$.
@@
There widely exist distribution shifts between data in different domains, i.e., $P^{d}(X,Y)\neq P^{d^{\prime}}(X,Y)$, where $d\neq d^{\prime}$. Inspired by (Zhang et al. 2013; Pearl, Glymour, and Jewell 2016), we use the causal graph shown in Figure 1 (a) to illustrate the causal model for data across domains. Specifically, as shown in Figure 1, we discuss the causal relations among five variables: label $Y$ (e.g., an object type), input features $X$ (e.g., an image), causal factors $C$ (e.g., the object shape in the image) that determine the label, non-causal factors $S$ (e.g., the background in the image), and domain variable $D$ (e.g., the data source). $C$ and $S$ are usually high-level latent concepts without observed supervision. Noticeably, $S$ often has spurious correlations with $Y$, even if they are not causally related. Such spurious correlations often vary in different domains, i.e., $P^{d}(Y|S)\neq P^{d^{\prime}}(Y|S)$.
@@
In the CMNIST dataset, it is unsurprising that a classifier relying on the digit color would fail on the test domain due to the shift in the spurious correlation $P(Y|S)$. To show the negative impact of spurious correlation on certified robustness, we compare the Average Certified Radius (ACR, a metric used to evaluate certified robustness) of random smoothing-based certified defense on CMNIST with the results on MNIST (where there is no digit color and thus the above spurious correlations do not exist). As observed from the results in Table 1, there is a significant degradation of prediction accuracy and ACR on the test domain, indicating severe issues for certified defense under domain shift.
@@
$||f(x_{1})-f(x_{2})||_{2}\leq{L}||x_{1}-x_{2}||_{2}.$ (4)
Based on the definition, if a neural network $f$ is 1-Lipschitz, then for any input $x$, the output $y$ satisfies $||\mathbf{y}||_{2}\leq||\mathbf{x}||_{2}$. Equivalently, $||\mathbf{y_{1}}-\mathbf{y_{2}}||_{2}\leq||\mathbf{x_{1}}-\mathbf{x_{2}}||_{2}$.
@@
where $\mathbb{C}$ and $\mathbb{S}$ are the space of $C$ and $S$, respectively. Here, $d\neq d^{\prime}$. Each of the equations above decomposes $P(Y|X)$ into two components. As shown in Eq. 6, the model can generalize to a different (or even adversarial) domain $d^{\prime}$ if it accurately captures the causal factors $C$ from $X$. The other term $P(Y|C)$ remains invariant across domains, which helps to mitigate the risk of increased vulnerability in new domains. However, as indicated by Eq. (7), $P^{d^{\prime}}(Y|X,s)$ varies across domains, which can increase the vulnerability to adversarial perturbations. This increased vulnerability stems from two main issues, as illustrated in Figure 1 (c): 1) reduced accuracy in the test domain, leading to diminished prediction reliability even under slight perturbations; and 2) the change in $P^{d^{\prime}}(Y|X,s)$ across domains increases decision uncertainty at each $X=x$ due to potential conflicts between $P(Y|C)$ and $P(Y|S)$. These factors collectively complicate the task of achieving robustness across different domains. The above analysis indicates the importance of incorporating a causal view into the robustness problem across domains. In our framework, we identify the causal factors from input (i.e., modeling $P(C|X)$) with certifiable robustness, and conduct certified defense based on an invariant predictor $P(Y|C)$.
Figure 2: An overview of the proposed framework GLEAN. The upper part represents the training process, while the lower part depicts the certification process on the test domain. Here, we showcase two training domains and one test domain with two classes 0 and 1, where the color of the object is a spurious factor. We define $S=0$ as orange and $S=1$ as green. In training domain 1, the spurious distribution between color and category is $P(Y=0|S=0)=0.9$ and $P(Y=1|S=1)=0.1$. These values change to 0.8 and 0.2, respectively in training domain 2, and then to 0.1 and 0.9 in the test domain. Thus, there is a correlation shift between the different domains of this dataset. The causal encoder is equipped with Lipschitz constraints with Lipschitz constant $L$. $Y_{train}$ and $Y_{test}$ are ground truth labels. $\hat{Y}$ is the predicted label. $z$ is the causal latent representations and each $\eta$ is a Gaussian noise. $y_{A}$ is the most probable class among all the $\hat{Y}$ after sampling with the probability $p_{A}$. Then we can leverage $p_{A}$ to compute the certified radius in latent space $CR_{z}$ and finally revert it back to get the certified radius $CR$ in input space.
@@
where $R^{d}(\beta\circ\Psi)=\mathbb{E}[\mathcal{L}(g(\Psi(x)),y)]$ is the prediction loss in domain $d$ with an encoder $\Psi$ and classifier $\beta$, $w$ is a “dummy” classifier and can be fixed as a scalar 1.0, and $\mathbb{D}_{\mathrm{tr}}$ is the set that includes all training domains. According to (Arjovsky et al. 2019), the gradient of $R^{d}(w(\beta\circ\Psi))$ reflects the invariance of the learned latent representations. The non-negative hyperparameter $\lambda$ controls the balance between the predictive ability and invariance.
@@
Certified Robustness for Unseen Domains
While significant progress has been made in certified defenses when training and test data share the same distribution, there is still limited exploration and a lack of theoretical guarantees for certified robustness under domain shifts. In this subsection, we bridge this gap by utilizing the theoretical support from certified defense (Cohen, Rosenfeld, and Kolter 2019) and causal inference to derive necessary theorems in this setting.
According to previous discussions, we perform random smoothing for the causal factors in the latent space. Therefore, based on the calculation of the certified radius, we introduce the following Theorem 1:
Theorem 1. Suppose we have a causal encoder $\Psi:\mathcal{X}\rightarrow\mathcal{Z}$, and an arbitrary classifier $\beta:\mathcal{Z}\rightarrow\mathcal{Y}$. Let $g_{\beta}$ be defined as $g_{\beta}(z)=\underset{y\in\mathcal{Y}}{\arg\max}P(\beta(z+\eta)=y)$, where $\eta\sim\mathcal{N}(0,\sigma^{2}I)$, $z=\Psi(x)$ is the latent causal representation. Suppose $\underline{p_{A}}$ is the lower bound of $p_{A}$, $\overline{p_{B}}$ is the upper bound of $p_{B}$ (here $\underline{p_{A}}$ and $\overline{p_{B}}$ are obtained based on $g_{\beta}$ with input in the representation space), $\underline{p_{A}},\overline{p_{B}}\in[0,1]$ and satisfy:
Then $\forall d,d^{\prime}\in\mathbb{D}$, $g_{\beta}(z^{d}+\delta_{z})=g_{\beta}(z^{d^{\prime}}+\delta_{z})=y_{A}$ for all $\|\delta_{z}\|_{2}<CR_{z}$, where $\delta_{z}$ is the perturbation applied to the latent causal representation $z$ and
Briefly, if we use an $L$-Lipschitz neural network in the causal factor learning module, we can calculate the certified radius in the input space. This is because we can simply scale the certified radius in the latent space by the Lipschitz constant $L$, such that $CR\geq{CR_{z}/L}$. If $L=1$, then $CR_{z}$ will be the lower bound of $CR$. With the aforementioned causal invariant assumption, the certified robustness for instances in one domain can also be propagated to instances in other domains with the same causal factors. Therefore, we are able to provide theoretical guarantees for cross-domain certified robustness. Detailed proofs can be found in the Appendix.
Implementation
Overview of Framework
Architecture
As aforementioned, we use Lipschitz constraints in the causal factor learning module. We define the final linear layer as the classifier $\beta$, with all preceding layers forming the encoder $\Psi$. We apply the Cayley transform (Trockman and Kolter 2021) to achieve orthogonality, thereby ensuring that each linear layer has a Lipschitz constant of 1. For the activation functions, we employ GroupSort (Anil, Lucas, and Grosse 2019), which also has 1-Lipschitzness. More details on the implementation of 1-Lipschitz networks can be found in the Appendix.
@@
  • Gaussian (Cohen, Rosenfeld, and Kolter 2019): Standard training with Gaussian noise based random smoothing.
@@
Evaluation Metrics. We consider two widely-used evaluation metrics: (1) certified accuracy at different radii, which is defined as the fraction of the test set that CERTIFY (Cohen, Rosenfeld, and Kolter 2019) classifies correctly. CERTIFY is a practical Monte Carlo-based certification procedure that offers the prediction of $g$ along with the lower bound of the certified radius or abstains the certification by sampling over $n$ Gaussian noises with the probability of at least $1-\alpha$, $\alpha$ is the significance level; (2) average certified radius (ACR), which is defined as $ACR=\frac{1}{|\mathbb{D}_{\mathtt{test}}|}\sum_{(x,y)\in\mathbb{D}_{\mathtt{test}}}{CR}(f;\sigma,x,y)\cdot\mathbf{1}_{[g(x,\sigma)=y]}$. Here, $|\mathbb{D}_{\mathtt{test}}|$ is the capacity of the test set, $CR$ is the certified radius returned by CERTIFY, $\mathbf{1}$ is the indicator function. We assign 0 to $CR$ for incorrect prediction of $g$. We use the same settings in (Cohen, Rosenfeld, and Kolter 2019) with $n=100000,n_{0}=100,\alpha=0.001$ to apply CERTIFY. Here $n_{0}$ is the small number of samples to find $y_{A}$. Note that, for two different models, their certified accuracies sometimes cannot be directly compared. At a specific radius $r$, one model may have a higher certified accuracy than the other, but the situation may be reversed at another radius. Therefore, ACR is a more suitable metric as it reflects “average robustness”.
Training Details. We use a three-layer MLP for CMNIST and a four-layer CNN for CelebA and DomainNet. During inference, we apply RS with the noise level $\sigma=0.12$. The result of other $\sigma$ is shown in the Appendix. We set the parameter of the regularization term $\lambda=10000$ for all datasets.

### v3

@@ Abstract
While machine learning models have proven effective across various scenarios, it is widely acknowledged that many models are vulnerable to adversarial attacks. Recently, there have emerged numerous efforts in adversarial defense. Among them, certified defense is well known for its theoretical guarantees against arbitrary adversarial perturbations on input within a certain range (e.g., $l_{2}$ ball). However, most existing works in this line struggle to generalize their certified robustness in other data domains with distribution shifts. This issue is rooted in the difficulty of eliminating the negative impact of spurious correlations on robustness in different domains. To address this problem, in this work, we propose a novel certified defense framework GLEAN, which incorporates a causal perspective into the generalization problem in certified defense. More specifically, our framework integrates a certifiable causal factor learning component to disentangle the causal relations and spurious correlations between input and label, and thereby exclude the negative effect of spurious correlations on defense. On top of that, we design a causally certified defense strategy to handle adversarial attacks on latent causal factors. In this way, our framework is not only robust against malicious noises on data in the training distribution but also can generalize its robustness across domains with distribution shifts. Extensive experiments on benchmark datasets validate the superiority of our framework in certified robustness generalization in different data domains. Code is available at https://github.com/yrqiao/Glean/tree/main.
@@ Introduction
Machine learning (ML) models, particularly deep neural networks (DNN), have demonstrated remarkable success across many areas (Devlin et al. 2018; Silver et al. 2017; He et al. 2016). Despite their success, these models still exhibit significant vulnerabilities to adversarial perturbations on input (Szegedy et al. 2013; Goodfellow et al. 2014; Biggio et al. 2013). A typical example in image classification is that a trained classifier that correctly classifies an image $x$ can be easily fooled by a perturbed image $x+\delta$, where $\delta$ represents adversarial perturbations imperceptible to human perceptions. This weakness impedes the deployment of ML models in critical applications where security and reliability are priorities, such as autonomous driving and healthcare.
@@
In the past few decades, researchers have developed numerous defense methods to enhance the adversarial robustness of ML models. Many of them are based on adversarial training (Goodfellow et al. 2014; Madry et al. 2017; Zhang et al. 2019; Athalye et al. 2018), which incorporates adversarial samples into model training. Despite its impressive performance, adversarial training is an empirical approach that lacks theoretical guarantees. That is, although it can enhance robustness against certain types of attacks, it may still be vulnerable to other unknown, or more potent adversarial perturbations. Differently, another line of work develops certified robustness. A certified robust classifier can theoretically guarantee that its prediction for a point $x$ remains constant within a certain specified range (a.k.a. radius) of perturbations on $x$, regardless of the type of attack. Randomized smoothing-based certified defense (Lecuyer et al. 2019; Li et al. 2018; Cohen et al. 2019) is one of the most representative methods in this area. Specifically, given an arbitrary base classifier $f$, this method can convert it to a certifiably robust classifier $g$, which is created by randomly sampling multiple noised versions of a given input and using the aggregated output from these variations to make final predictions. Inspired by this approach, many subsequent studies (Li et al. 2019; Jeong and Shin 2020; Jeong et al. 2021; Salman et al. 2019; Zhai et al. 2020) have expanded upon the basis of random smoothing.
@@
Currently, most existing certified defense works focus on data in the same domain yet overlook other domains with distribution shifts. This limitation can result in a markedly degraded certified robustness performance when these methods are applied to the test domain (Sun et al. 2021). As discussed in previous work (Ilyas et al. 2019; Beery et al. 2018), such degradation of robustness lies in the fact that ML models tend to overfit spurious correlations between features and labels. As these spurious correlations often vary across different domains (Ye et al. 2022), fitting spurious correlations can easily lead $g$ to make incorrect predictions or correct predictions but with lower confidence levels. The former results in the certified radius being assigned zero, while the latter also leads to a reduced certified radius. Therefore, domain shifts can lead to weak generalization w.r.t. not only prediction performance but also certified robustness. Different from ML models, humans can naturally capture the invariant relations between labels and their corresponding causal factors, various studies (Zhang et al. 2020; Tang et al. 2020; Schölkopf et al. 2021) argue that human’s inherent causal view brings a solution to avoid the domain generalization hurdle for robustness. Inspired by this, in this paper, we study the problem of generalizing the certified robustness under domain shifts from a causal view.
@@
However, addressing this problem presents multifaceted challenges. Challenge 1: As aforementioned, spurious correlations varying across domains adversely affect robustness. To achieve robustness across domains, it is crucial to effectively remove the impact of these spurious correlations. However, identifying and eliminating the impact of spurious correlations on robustness in different domains presents a challenge. Challenge 2: Apart from spurious factors, the distribution shifts often bring challenges for the model to defend against the perturbations on the factors that causally determine the label in unseen domains, which leads to diminished certified robustness. Challenge 3: It is important to provide theoretical guarantees for robustness on other data domains, but most existing works remain in empirical observations and lack theoretical analysis. Although some work in certified defenses has provided upper bounds on perturbations while maintaining robustness, they were not designed to address certified robustness in the domain shift context.
@@
In this work, to tackle these challenges, we propose a novel framework GeneraLizable cErtified cAusal defeNse (GLEAN) that enhances the certified robustness of models on data in different domains. To mitigate the influence of spurious correlations on robustness generalization (Challenge 1), we construct a causal model for data in different domains, and simultaneously conduct a causal analysis on model robustness and generalization. Based on the causal model, we filter out the impact from spurious correlations and enhance robustness across domains. This is different from most existing defense algorithms which take the same strategy indiscriminately towards all the input features. To achieve certified robustness through causal factors (Challenges 2), we utilize a certified causal factor learning module with Lipschitz constraint. This module enables certification through the latent representation space for high-level causal factors, conducting certified defense for perturbations on causal factors that determine the label in different domains. To bring theoretical guarantees for robustness on different data domains (Challenge 3), we derive a theoretical analysis by leveraging the theoretical support of certified defense and causal inference. Our main contributions can be summarized as:
@@
Many previous works proposed certification methods to derive a certified radius that is the lower bound of the robust radius. Researches in this area fall into two categories: exact methods and conservative methods. Exact methods (Ehlers 2017; Bunel et al. 2018; Tjeng et al. 2017), usually based on Satisfiability Modulo Theories or mixed integer linear programming, guarantee the identification of a perturbation $\delta$ within a radius $r$ that can cause $f$ to change its prediction. However, they require the model to have a limited scale. Conservative methods (Wong and Kolter 2018; Wong et al. 2018; Gowal et al. 2018) ensure the detection of existing adversarial examples and, in addition, refuse to make certification for some vulnerable data points. These methods, though more scalable, impose specific assumptions on the model’s architecture.
@@
Randomized Smoothing (RS) is proposed to tackle the above limitations, which can be applied to any architectures (Cohen et al. 2019). It constructs a smoothed classifier $g$ from an arbitrary base classifier $f$. The definition of $g$ is as follows:
@@
In this formula, $\eta\sim\mathcal{N}(0,\sigma^{2}\bm{I})$ is the isotropic Gaussian noise with the noise level $\sigma$ as the hyperparameter. The smoothed classifier $g$ can be summarized as returning the class most likely to be predicted by $f$ when the input $x$ is sampled from Gaussian distributions. (Cohen et al. 2019) provide the theoretical form of certified radius which is the lower bound of the robust radius:
@@
where $p_{A}=P(f(x+\eta)=y_{A})$, $p_{B}=\max_{y\neq y_{A}}P(f(x+\eta)=y)$, meaning that $f$ will mostly return class $y_{A}$ with the probability $p_{A}$, and will return the runner-up class with the probability $p_{B}$. $\Phi^{-1}$ is the inverse of the standard Gaussian cumulative distribution function. Then $g(x+\delta)=y_{A}$ for all $\delta||_{2}\leq{CR}$.
@@
There widely exist distribution shifts between data in different domains, i.e., $P^{d}(X,Y)\neq P^{d^{\prime}}(X,Y)$, where $d\neq d^{\prime}$. Inspired by (Zhang et al. 2013; Pearl et al. 2016), we use the causal graph shown in Figure 1 (a) to illustrate the causal model for data across domains. Specifically, as shown in Figure 1, we discuss the causal relations among five variables: label $Y$ (e.g., an object type), input features $X$ (e.g., an image), causal factors $C$ (e.g., the object shape in the image) that determine the label, non-causal factors $S$ (e.g., the background in the image), and domain variable $D$ (e.g., the data source). $C$ and $S$ are usually high-level latent concepts without observed supervision. Noticeably, $S$ often has spurious correlations with $Y$, even if they are not causally related. Such spurious correlations often vary in different domains, i.e., $P^{d}(Y|S)\neq P^{d^{\prime}}(Y|S)$.
@@
In the CMNIST dataset, it is unsurprising that a classifier relying on the digit color would fail on the test domain due to the shift in the spurious correlation $P(Y|S)$. To show the negative impact of spurious correlation on certified robustness, we compare the Average Certified Radius (ACR, a metric used to evaulate certified robustness) of random smoothing-based certified defense on CMNIST with the results on MNIST (where there is no digit color and thus the above spurious correlations do not exist). As observed from the results in Table 1, there is a significant degradation of prediction accuracy and ACR on the test domain, indicating severe issues for certified defense under domain shift.
@@
[…]
@@
where $\mathbb{C}$ and $\mathbb{S}$ are the space of $C$ and $S$, respectively. Here, $d\neq d^{\prime}$. Each of the equations above decomposes $P(Y|X)$ into two components. As shown in Eq. 6, the model can generalize to a different (or even adversarial) domain $d^{\prime}$ if it accurately captures the causal factors $C$ from $X$. The other term $P(Y|C)$ remains invariant across domains, which helps to mitigate the risk of increased vulnerability in new domains. However, as indicated by Eq. (7), $P^{d^{\prime}}(Y|X,s)$ varies across domains, which can increase the vulnerability to adversarial perturbations. This increased vulnerability stems from two main issues, as illustrated in Figure 1 (c): 1) reduced accuracy in the test domain, leading to diminished prediction reliability even under slight perturbations; and 2) the change in $P^{d^{\prime}}(Y|X,s)$ across domains increases decision uncertainty at each $X=x$ due to potential conflicts between $P(Y|C)$ and $P(Y|S)$. These factors collectively complicate the task of achieving robustness across different domains. The above analysis indicates the importance of incorporating a causal view into the robustness problem across domains. In our framework, we identify the causal factors from input (i.e., modeling $P(C|X)$) with certifable robustness, and conduct certified defense based on an invariant predictor $P(Y|C)$.
Figure 2: An overview of the proposed framework GLEAN. The upper part represents the training process, while the lower part depicts the certification process on test domain. Here, we showcase two training domains and one test domain with two classes 0 and 1, where the color of the object is a spurious factor. We define $S=0$ as orange and $S=1$ as green. In training domain 1, the spurious distribution between color and category is $P(Y=0|S=0)=0.9$ and $P(Y=1|S=1)=0.1$. These values change to 0.8 and 0.2, respectively in training domain 2, and then to 0.1 and 0.9 in the test domain. Thus, there is a correlation shift between the different domains of this dataset. The causal encoder is equipped with Lipschitz constraints with Lipschitz constant $L$. $Y_{train}$ and $Y_{test}$ are ground truth labels. $\hat{Y}$ is the predicted label. $z$ is the causal latent representations and each $\eta$ is a Gaussian noise. $y_{A}$ is the most probable class among all the $\hat{Y}$ after sampling with the probability $p_{A}$. Then we can leverage $p_{A}$ to compute the certified radius in latent space $CR_{z}$, and finally revert it back to get the certified radius $CR$ in input space.
@@
where $R^{d}(\beta\circ\Psi)=\mathbb{E}[\mathcal{L}(g(\Psi(x)),y)]$ is the prediction loss in domain $d$ with an encoder $\Psi$ and classifier $\beta$. $w$ is a “dummy” classifier and can be fixed as a scalar 1.0. According to (Arjovsky et al. 2019), the gradient of $R^{d}(w(\beta\circ\Psi))$ reflects the invariance of the learned latent representations. The non-negative hyperparameter $\lambda$ controls the balance between the predictive ability and invariance.
@@
Certified Robustness for Unseen Domains
While significant progress has been made in certified defenses when training and test data share the same distribution, there is still limited exploration and a lack of theoretical guarantees for certified robustness under domain shifts. In this subsection, we bridge this gap by utilizing the theoretical support from certified defense (Cohen et al. 2019) and causal inference to derive necessary theorems in this setting. According to previous discussions, we perform random smoothing for the causal factors in the latent space. Therefore, based on the calculation of the certified radius, we introduce the following Theorem 1:
Theorem 1. Suppose we have an causal encoder $\Psi:\mathcal{X}\rightarrow\mathcal{Z}$, and an arbitrary classifier $\beta:\mathcal{Z}\rightarrow\mathcal{Y}$. Let $g_{\beta}$ be defined as $g_{\beta}(z)=\underset{y\in\mathcal{Y}}{\arg\max}P(\beta(z+\eta)=y)$, where $\eta\sim\mathcal{N}(0,\sigma^{2}I)$, $z=\Psi(x)$ is the latent causal representation. Suppose $\underline{p_{A}}$ is the lower bound of $p_{A}$, $\overline{p_{B}}$ is the upper bound of $p_{B}$, $\underline{p_{A}},\overline{p_{B}}\in[0,1]$ and satisfy:
Then $\forall d,d^{\prime}\in\mathbb{D}$, $g_{\beta}(z^{d}+\delta_{z})=g_{\beta}(z^{d^{\prime}}+\delta_{z})=y_{A}$ for all $\|\delta_{z}\|_{2}<CR_{z}$, where $\delta_{z}$ is the perturbation applied to latent causal representation $z$ and
Briefly, if we use an $L$-Lipschitz neural network in the causal factor learning module, we can calculate the certified radius in the input space. This is because we can simply scale the certified radius in the latent space by the Lipschitz constant $L$, such that $CR\geq{CR_{z}/L}$. If $L=1$, then $CR_{z}$ will be the lower bound of $CR$. With the aforementioned causal invariant assumption, the certified robustness for instances in one domain can also be propagated to instances in other domains with the same causal factors. Therefore, we are able to provide theoretical guarantees for cross-domain certified robustness. Detailed proofs of Theorem 1 and Theorem 2 can be found in the Appendix.
Implementation
Overview of Framework
Architecture
As aforementioned, we use Lipschitz constraints in the causal factor learning module. We define the final linear layer as the classifier $\beta$, with all preceding layers forming the encoder $\Psi$. We apply the Cayley transform (Trockman and Kolter 2021) to achieve orthogonality, thereby ensuring that each linear layer has a Lipschitz constant of 1. For the activation functions, we employ GroupSort (Anil et al. 2019), which also has 1-Lipschitzness. More details on the implementation of 1-Lipschitz networks can be found in the Appendix.
@@
  • Gaussian (Cohen et al. 2019): Standard training with Gaussian noise based random smoothing.
@@
Evaluation Metrics. We consider two widely-used evaluation metrics: (1) certified accuracy at different radii, which is defined as the fraction of the test set that CERTIFY (Cohen et al. 2019) classifies correctly. CERTIFY is a practical Monte Carlo-based certification procedure that offers the prediction of $g$ along with the lower bound of the certified radius or abstains the certification by sampling over $n$ Gaussian noises with the probability of at least $1-\alpha$, $\alpha$ is the significance level; (2) average certified radius (ACR), which is defined as $ACR=\frac{1}{|\mathbb{D}_{\mathtt{test}}|}\sum_{(x,y)\in\mathbb{D}_{\mathtt{test}}}{CR}(f;\sigma,x,y)\cdot\mathbf{1}_{[g(x,\sigma)=y]}$. Here, $|\mathbb{D}_{\mathtt{test}}|$ is the capacity of the test set, $CR$ is the certified radius returned by CERTIFY, $\mathbf{1}$ is the indicator function. We assign 0 to $CR$ for incorrect prediction of $g$. We use the same settings in (Cohen et al. 2019) with $n=100000,n_{0}=100,\alpha=0.001$ to apply CERFITY. Here $n_{0}$ is the small number of samples to find $y_{A}$. Note that, for two different models, their certified accuracies sometimes cannot be directly compared. At a specific radius $r$, one model may have a higher certified accuracy than the other, but the situation may be reversed at another radius. Therefore, ACR is a more suitable metric as it reflects “average robustness”.
Training Details. We use a three-layer MLP for CMNIST and a four-layer CNN for CelebA and DomainNet. During inference, we apply RS with the noise level $\sigma=0.12$. The result of other $\sigma$ is shown in Appendix. We set the parameter of the regularization term $\lambda=10000$ for all datasets.